Could someone help me solve the substitution method using this example.
2x + 7y =12
x=8 -4y
What is the soltuion

Answer by stanbon(48535)

2x + 7y =12
x=8 -4y
-----------
Substitute x= 8-4y into the 1st equation:
2(8-4y) + 7y = 12
16-8y + 7y = 12
16 -y = 12
y = 4
---------------
Substitute into x = 8-4y to solve for Y.
x = 8-4*4
x = -8
------------
Solution
x = -8; y = 4
===================
Cheers,
Stan H.
Cheers,